OpenStudy (anonymous):

Equations with fractions; Show work 15m + 3/2 = -6

OpenStudy (anonymous):

I know I would multiply both sides by 2 right?

OpenStudy (anonymous):

You have to multiply EACH TERM on both sides by 2

OpenStudy (anonymous):

Then would I subtract 3 from both sides or is there something I have to do before that!

OpenStudy (anonymous):

correct you have to subtract 3 from both sides

OpenStudy (austinl):

15m + 3/2 = -6 15m = -15/2 m = -1/2

OpenStudy (austinl):

If you are unafraid of fractions. :D

OpenStudy (anonymous):

That is not what I got! I got 15m + 3 = -12 15m + 3 -3 = -12 - 3 15m/15 = -15/15 m= -1 Is that wrong!

OpenStudy (anonymous):

Hmmm, I have gotten 3 different answers now!

OpenStudy (austinl):

15m + 3/2 = -6 30m+3 = -12 30m = -15 m = -1/2 Two different ways, same answer.

OpenStudy (anonymous):

15m + 3/2 = -6 2(15m + 3/2) = 2(-6) 30m + 3 = -12 30m = -12-3 m = -15/30 m = -1/2 THIS IS THE RIGHT ANSWER!!!

OpenStudy (austinl):

As well, I verified with my calculator. And if you plug it back in, it works.

OpenStudy (anonymous):

skullpatrol started with the wrong initial equation!!!

OpenStudy (anonymous):

You guys do know that 15m + 3 is over 2 right! Not 3/2. I someone put it as 3/2.

OpenStudy (anonymous):

Then you should have written it as (15m+3)/2...

OpenStudy (austinl):

\[\frac{15m+3}{2} = -6\] USE PROPER FORMAT!!! That is the single most aggravating thing you can do.....

OpenStudy (austinl):

(15m+3)/2 = -6 15m + 3 = -12 15m = -15 m = -1

OpenStudy (anonymous):

(15m + 3)/2 = -6 2(15m + 3)/2 = 2(-6) 15m + 3 = -12 15m = -12-3 m = -15/15 m = -1

OpenStudy (austinl):

This illustrates EXACTLY why you should use proper format whilst you are asking and answering a question. It clears up confusion and expedites assistance to the asker.